OpenStudy (anonymous):

HELP ME PLEASE!! A geo sequence t7=192, t12=192 t3= ??

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

@cupcakezz the 7th and 12th terms are the same?

OpenStudy (anonymous):

Maybe the common ratio is 1?

OpenStudy (anonymous):

OH SORRY T12=6144 @genius12

OpenStudy (anonymous):

mmk A geometric sequence is in the form a, ar, ar^2...ar^(n-1). The seventh term in the given sequence is 192 which we can write as ar^6 = 192. The 12th term is 6144 which we can write as ar^(11) = 6144. Now if we divide both sides of these equations we get: ar^(11) 6144 ------ = ----- ar^6 192 The a's cancel out and this simplifies to: r^5 = 32 Taking the fifth root of both sides we get: r = (32)^(1/5) <-- note that I'm using 1/5 to denote the fifth root of 32. Now we know the common ratio. We now need to solve for 'a'. We can simply plug back in the value of 'r' that we just got in to one of the previous equations we used to find the value of 'a'. I'll plug in 'r' in to t7: t7 = 192 = ar^6 = a[(32)^(1/5)]^6 Which gives: a[(32)^(1/5)]^6 = 192 192 => a = ------------- = 3 [(32)^(1/5)]^6 Now we know both the values of 'a' and 'r'. Now we can find t3. Well t3 is just ar^(3-1) = ar^2. We know a = 3 and r = (32)^(1/5). Just plug them in and find the value :D. @cupcakezz

OpenStudy (anonymous):

I got the answer so when you get your answer, tell me and we can check.

OpenStudy (anonymous):

Btw I forgot to mention that 32 = 2^5 so we can re-write r = (32)^(1/5) as: r = (32)^(1/5) = (2^5)^(1/5) = 2. Hence r = 2 and we deduced a = 3. Now we have nice looking whole numbers.

OpenStudy (anonymous):

did you get 28?

OpenStudy (anonymous):

nope.

OpenStudy (anonymous):

Remember, a = 3 and r = 2 and t3 = ar^2. Just plug the numbers in..

OpenStudy (anonymous):

12!!

OpenStudy (anonymous):

can you help me with this one @genius12

OpenStudy (anonymous):

It could be that the sequence is \(ar^n\), and not \(ar^{n-1}\). Or anything else, for that matter. @genius12

OpenStudy (anonymous):

@genius12

OpenStudy (anonymous):

@SithsAndGiggles wat?

OpenStudy (anonymous):

why doesnt sin(x)=2 have no solution? @genius12

OpenStudy (anonymous):

@zzr0ck3r ^^ help

OpenStudy (anonymous):

sin(x) has a range of -1 <= y <= 1. It never goes to any number greater than 1 or any number lower than -1.